I have 2 problems, hear they are.
A gallon of vinegar is labelled 5% acidity. How many fluid Ounces of pure acid must be added to get 6% acidity?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

A gallon of vinegar is labled 5% acidity. How many fluid Ounces of pure acid must be added to get 6% acidity?
--------------------------
acid + acid = acid
0.5*1 + 1*x = 0.6(1+x)
0.5 + x = 0.6 + 0.6x
0.4x = 0.1
x = 1/4 gallon (amount of pure acid needed in the mixture)
======================

A 12% saline solution is to be deluted with water to obtain 20 oz. of a 10% solution. How much water is needed?
saline + saline = saline
0.12x + 0*(20-x)= 0.10*20
0.12x = 2
x = 16 2/3
water = 20-16 2/3 = 3 1/3 oz (amount of water needed for the mixture)
========================================================================
Cheers,
Stan H.

Answer by richwmiller(17219)

Guess what? This website has rules. "Hear" are some of them.
One problem at a time.
No similar problems.
Limit 4 problems daily.
(128*.05)+x=.06*(x+128)
x=1.3617